# Insight-HXMT View of the BHC Swift J1727.8-1613 during its outburst in 2023

### v1

@@ 4 Discussions
The Galactic black hole Swift J1727.8-1613 was monitored daily by various X-ray satellites like NICER, NuSTAR, Insight-HXMT, etc. Multi-wavelength follow-up observations were also carried out for this source. We have taken HXMT data for both our spectral and timing analysis from MJD 60181.4 to 60198.0. The evolution of the light curve makes the outburst nature of the source a fast-rise slow decay type. After the onset of the outburst, it rose to its peak very fast within 5 days. Using the $0.01~{}s$ time-binned light curves from the three instruments of HXMT (LE, ME, and HE), we studied the source’s timing properties by analyzing the light curves. Using the spectra files from these three instruments we also studied the spectral properties of this source by analyzing combined LE+ME+HE spectra in the $2-150$ keV broad energy range.
@@ 4.1 Propagation of Shock
Quasi-periodic oscillation is one of the most important and common phenomena for stellar mass black holes. For this newly discovered source, we analyzed light curves of a total of 342 light curves (114 for LE, ME, and HE each). However, in some of the exposures, light curves were not produced properly. There were a total of 92 exposures which commonly produced LE, ME, and HE light curves. We list those data in Table 3. Our analysis is final for 276 data. We found the presence of QPO in all of these light curves and they show rapid evolution in frequency in this short analysis period. The QPO frequency ($\nu_{qpo}$) showed an increase in frequency within even a single day. This is one of the advantages of using Insight-HXMT as it provides multiple exposures for one observation ID which helps to study the short-scale variations of such important physical properties of the source. The evolution of the QPO frequency has been discussed broadly in the result section. It showed an increasing trend from the start of our analysis. From the variation of QPO frequency, (%)RMS, and $Q$-factor, the nature of the QPO is designated to be type-C. Although this is a well-established topic in the literature, the origin of QPO is still a matter of debate. Many models try to explain the type-C QPO phenomenon, e.g., the Lense-Thirring effect (Stella & Vietri 1998), accretion-ejection instability (Tagger & Pellat 1999), oscillation of the disk (Titarchuk & Osherovich 2000), the misalignment between the spin and binary rotational axis (Ingram et al. 2009, and references therein).
@@
We would like to focus on a physical scenario that the QPOs could have originated from the shock instabilities in advective flows around black holes. Black hole accretion is transonic in which multiple sonic points could exist (Chakrabarti & Titarchuk 1995). Depending on the flow properties, there is the possibility of the formation of standing shocks (Singh et al. 2022 and references therein). The matter from the companion need not necessarily be only Keplerian. There could be a supply of matter which has angular momentum distribution deviated from the Keplerian one. This is the sub-Keplerian component. Such a lower angular momentum component of infalling matter gets accreted in a free-fall timescale. Due to the tug of war between the pressure of gravitational force and centrifugal force, this matter could virtually stop from falling onto the BH at some distance from the BH and undergo shock transition, creating a post-shock region. This approach could very well describe the spectral and temporal properties of BH, as it has been shown previously in both theoretical (Chakrabarti & Titarchuk 1995) and observational (Debnath et al. 2014; Mondal et al. 2014, Chatterjee et al. 2020, 2021, 2023) studies. The change in size of such post-shock region can describe the spectro-temporal properties during an outburst. According to Molteni et al. (1996), the oscillation in the shock region could have originated due to the presence of cooling. This was later shown in simulation works by Garain et al. (2014). The oscillation of the shock region could have originated due to the resonance condition between the cooling timescale due to the Comptonization process and the infall timescale in the coronal region (see Chakrabarti et al. 2015). Additionally, it has also been observed by the authors that once the QPO sets in, it gets locked for a week or more depending on the above condition. From the expressions of infall and cooling timescales, we get that the QPO frequency ($\nu_{qpo}$) could be given as (Molteni et al. 1996; Chakrabarti & Manickam 2000; Chakrabarti et al. 2004),
$\nu_{qpo}=\frac{c^{3}}{2GM_{BH}}\frac{1}{RX_{s}(X_{s}-1)^{1/2}}Hz,$ (1)
@@
where $c$ and $G$ are the speed of light and gravitational constant respectively. $M_{BH}$ represents the mass of the black hole, $X_{s}$ is the distance of the shock, or called shock location. $X_{s}$ is in the units of Schwarzschild radius ($r_{s}=2GM_{BH}/c^{2}$). $R$ is the ratio of matter densities in post-shock to pre-shock regions ($\rho_{+}/\rho_{-}$) and is known as the compression ratio. From our timing analysis, we already extracted the information about QPO frequency ($\nu_{qpo}$). Using the above relation, we estimated the shock location during the outburst. According to Chakrabarti et al. (2005), $X_{s}$ could be located anywhere from 10 to 1000 $r_{s}$. The shock stays very far away $\sim 1000~{}r_{s}$ when the spectral nature of the ongoing outburst is hard and becomes smaller gradually in the process of cooling. Using the relation (1), we found that the shock was far at the start of the analysis period (Fig. 8b). It then rapidly decreased as the QPO frequency increased rapidly during the initial few days. Close to MJD 60185, when the $\nu_{qpo}$ stopped its rapid increment, and the shock location started to decrease very slowly. The values of the shock location along with QPO frequency are given in Table 3. As equation (1) suggests, as the shock was bigger at the start of the outburst, the QPO frequency was small. As it decreased rapidly due to cooling, the $\nu_{qpo}$ increased rapidly. Then when the cooling was less efficient, the size of the shock changed very slowly, although decreasing. It made the $\nu_{qpo}$ increase with time but not so rapidly as before.
@@
Using this relation, we estimated the velocity of shock. This is given in Fig. 8c. We find that the shock velocity has varied within this period. In the beginning, it attained higher values when the $X_{s}$ was also larger, however decreasing rapidly. It slowed down as the shock moved inwards, along with some fluctuations. This could be due to the limitation of assuming constant velocity of the movement of the shock.
@@ 4.2 Evolution of the physical configuration and spectral nature
From the spectral properties, we have seen that the $T_{in}$ was first low and it gradually increased. The exact opposite trend was observed for the Normalization of the diskbb model. The norm of diskbb model varies as
@@
where, $R_{in}$, $D_{10}$ are the inner-disk radius of the accretion disk and the distance of the source in $10$ kpc unit. $i$ is the inclination angle of the accretion disk to the observer. Considering, $D_{10}$ and $\theta$ do not change during an outburst, the relation could be shortened like $R_{in}\sim N_{diskbb}^{1/2}$. Thus as the $N_{diskbb}$ decreased, it means the inner edge of the disk moved closer to the black hole. This is also supported by the fact that $T_{in}$ increased as the disk came closer.
This also supports the properties we showed from the timing analysis. The shock forms at a distance ($X_{s}$) when the sub-Keplerian matter is halted due to the counterbalance between the pressures of gravitational force and the centrifugal force. This forms the hot Compton cloud, as it becomes hot and puffed up in time. The hot Compton cloud will intercept some of those disk photons and there will be scattering in the process of inverse-Comptonization that will increase the energy of those intercepted photons and will emit them at higher energies. In this way, the Compton cloud will be cooled due to energy loss. Due to this cooling, the Compton cloud will gradually shrink in size. As it continues shrinking in size, the disk will continue to move inwards. As we have seen from the timing properties, the shock location was larger at the start of the outburst. The inner-disk radius was also higher as $R_{in}\sim N_{diskbb}^{1/2}$ and $N_{diskbb}$ was higher. Then cooling starts as the disk starts to approach inwards. In this way, the Compton cloud was losing energy by up-scattering and as the cooling continued, the $X_{s}$ and $R_{in}$ were both moving closer to the black hole.
Looking at the spectral parameters, and also the location of the shock, we designate that the source was already in an intermediate state at the start of our analysis period. As the outburst progressed, it was transitioning towards softer states.
From the values of the parameter $\cos(i)$ of the pexrav model, we find that the inclination of the disk was $>60^{\circ}$, which makes this source a high inclination binary system.
The hydrogen column density ($N_{H}$) showed a variation, which is common for Galactic black hole binaries unless the system is heavily obscured, which has recently been found for the newly discovered BH Swift J151857.0-572147 (Mondal et al. 2024). The average $N_{H}$ has an average value of $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
@@ 5 Summary and Conclusions
(vi) This source is in a high inclination binary system.

### v2

@@ 4 Discussions
The Galactic black hole Swift J1727.8-1613 went through an outburst recently in 2023. We have used Insight-HXMT data for our both spectral and timing studies from MJD 60181.4 to 60198.0. The evolution of the light curve makes the outburst nature of the source a fast-rise slow decay type. After the onset of the outburst, the flux reached its peak value very fast within 5 days. Using the $0.01$ sec time-binned light curves from the three instruments of HXMT (LE, ME, and HE), we studied the source’s timing properties. Using the spectra files from these three instruments, we also studied the spectral properties of this source by analyzing combined LE+ME+HE spectra in the $2-150$ keV broad energy range.
@@ 4.1 Propagation of Shock
Quasi-periodic oscillation is one of the most important and common phenomena for stellar mass black holes. For this newly discovered source, we analyzed a total of 342 light curves (114 for LE, ME, and HE each). However, in some of the exposures, light curves were not produced properly. There were a total of 92 exposures which commonly produced LE, ME, and HE light curves. We list those data in Table 3. Finally, we analyzed 276 lightcurves. We found the presence of QPO in all of these light curves, and they showed rapid evolution in frequency in this short analysis period. The QPO frequency ($\nu_{qpo}$) showed an increase in frequency within even a single day. The evolution of the QPO frequency has been discussed broadly in the result section. It showed an increasing trend from the start of our analysis. From the variation of QPO frequency, (%)RMS, and $Q$-factor, the nature of the QPO is designated to be type-C. Although the QPOs are extensively studied in the literature, their origin requires in depth modeling.
@@
Here, we would like to focus on the physical scenario that explains the origin of QPOs due to the shock instabilities in advective flows around BHs. Accretion onto BH is transonic flow with possibility of multiple sonic points (Chakrabarti 1989). The matter from the companion need not necessarily be only Keplerian. There could be a supply of matter which has angular momentum distribution deviated from the Keplerian one. This is the sub-Keplerian component. Such a lower angular momentum component of infalling matter gets accreted in a free-fall timescale. Due to the tug of war between the gravitational force and centrifugal force, this matter could virtually stop at some distance from the BH and undergo shock transition, creating a post-shock region. Depending on the flow properties, the standing shocks forms (Chakrabarti 1989; Singh et al. 2022 and references therein). This approach can well describe the observed spectral and temporal properties of BHs (Debnath et al. 2014; Mondal et al. 2014, Chatterjee et al. 2020, 2021, 2023). According to Molteni et al. (1996), the shock can be oscillatory due to the presence of cooling. This was later shown in simulation by Garain et al. (2014). During the oscillation when the cooling timescale due to the Comptonization process and the heating timescale matches, QPOs originate (see Chakrabarti et al. 2015). Additionally, it has also been observed by the authors that once the QPO sets in, it gets locked for a week or more depending on the above condition. From the expressions of infall and cooling timescales, we get that the QPO frequency ($\nu_{qpo}$) can be estimated using the relation (Molteni et al. 1996; Chakrabarti & Manickam 2000; Chakrabarti et al. 2004),
@@
Where $c$, $G$, $M_{BH}$, $X_{s}$, and $R$ are the speed of light, gravitational constant, the mass of the BH, location of the shock, and the ratio of matter densities in post-shock to pre-shock regions ($\rho_{+}/\rho_{-}$) respectively. From our timing analysis, we already extracted the information about QPO frequency ($\nu_{qpo}$). Using the above relation, we estimated the shock location during the outburst. According to Chakrabarti et al. (2005), $X_{s}$ can be located anywhere above 10 $r_{s}$ depending on the flow parameters producing shocks. The shock forms far away $\sim 1000~{}r_{s}$ when the spectral nature of the ongoing outburst is hard and decreases gradually in the progressive days as cooling increases (Mondal et al. 2015). Using Eq. 1, we found that the shock was far away from the BH at the start of the outburst (see Figure 8b). It then significantly decreased as the QPO frequency increased rapidly during the initial few days. Close to MJD 60185, when the $\nu_{qpo}$ stopped increasing, and the shock location started to decrease very slowly. The values of the shock location along with QPO frequency are given in Table 3.
@@
Using this relation, we have estimated the velocity of the movement of the shock. The estimated shock velocities are given in Figure 8c, which is a few tens of m/s in accord with the previous estimations (Chakrabarti et al. 2005; Mondal et al. 2015). We find that the shock velocity has varied within this period. At the beginning, it attained higher values when the $X_{s}$ was also larger, however decreasing significantly afterward. It slowed down as the shock moved inwards, along with some fluctuations. This can be due to the limitation of assuming constant velocity of the shock propagation.
@@ 4.2 Evolution of the physical configuration and spectral nature
From the spectral properties, we have seen that the $T_{in}$ was low initially and then gradually increased. The exact opposite trend was observed for the Normalization of the diskbb model. The norm of diskbb model varies as
@@
where, $R_{in}$, $D_{10}$ are the inner-disk radius of the accretion disk and the distance of the source in $10$ kpc unit. $i$ is the inclination angle of the accretion disk to the observer. Assuming that $D_{10}$ and $\theta$ do not change during an outburst, the relation could be shortened like $R_{in}\sim N_{diskbb}^{1/2}$. For our case, except for one exposure, the inclination varied in a narrow range of 78-86. Thus, as the $N_{diskbb}$ decreased, it means the inner edge of the disk moved closer to the BH, which agrees with the fact that $T_{in}$ increased as the disk moved closer. Furthermore, this also supports the properties we showed from the timing analysis that the shrinking of post-shock region with time.
Looking at the spectral parameters, and also the location of the shock, we designate that the source was already in an intermediate state at the start of our analysis period. As the outburst progressed, it was transitioning towards softer states. From the values of the parameter $\cos(i)$ of the pexrav model, we find that the inclination of the disk was $>60^{\circ}$, which makes this source a highly inclined system. The hydrogen column density ($N_{H}$) showed a variation, which is common for Galactic black hole binaries unless the system is heavily obscured, which has recently been found for the newly discovered BH Swift J151857.0-572147 (Mondal et al. 2024). The average $N_{H}$ has an average value of $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
@@ 5 Summary and Conclusions
(vi) This source has a high inclination.